OpenStudy (anonymous):

what is 2+9 over 4=2.1

OpenStudy (diyadiya):

11/4=2.75

OpenStudy (anonymous):

11/4=2.1 not true

jhonyy9 (jhonyy9):

2+9/4 =(8+9)/4 =17/4

OpenStudy (anonymous):

2.75